One mph is equal to approximately 0.869 knots. What is the average speed in knots of a car that makes a 90 mile trip in 5.5 hours?
Answer by TimothyLamb(4379)

s = d/t
s = 90/5.5 mi/hr
---
s = 90/5.5 mi/hr * (0.868976 nm/mi)
s = 14.2196072727
---
answer:
14.22 knots
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---